# Existence of cohomogeneity one Einstein metrics

### v4

@@ 4. Global Analysis
Motivated by the derivative formula (4.3) and Proposition 4.3, we study the sign of $Q$ when the function is restricted on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. The set $\mathfrak{S}_{A}(\sigma,\tau)$ is equivalently defined as
@@
For any $d_{2}\geq d_{1}\geq 2$, the inequality $P_{X}\geq 0$ holds on $[0,1]$ and only vanishes at $k=1$.
@@
The coefficient of each $k^{i}$ is positive on $[0,1]$ since $d_{1},d_{2}\geq 2$. Therefore, the polynomial is positive on $[0,1]$. ∎
@@
The critical point $p_{0}^{+}$, whose $(k,l)$-coordinate is $(0,0)$, is on the boundary of $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. If $A=0$, the product Einstein metric is represented by the algebraic curve $\mathcal{E}\cap\{X_{2}=0\}\cap\left\{R_{2}-\frac{1-H^{2}}{n}=0\right\}$, at which both $P$ and $Q$ vanish. This is reflected by $\omega(0,0,l)$ being identically zero.
For any $A\geq 0$, both $\omega(A,k,l)$ and its gradient vanish at $(0,0)$. To learn the sign of $\omega(A,k,l)$ near $p_{0}^{+}$, we compute the Hessian of the function at $(0,0)$.
@@
By the second partial test, we have the following proposition.
@@
Note that $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$ is smaller than the upper bound in Proposition 4.3 if $d_{2}\geq d_{1}\geq 2$. By the derivative formula (4.3), Proposition 4.3 and Proposition 4.5 indicate that if any integral curve in $\mathfrak{S}_{A}(\sigma,\tau)$ does not escape the set only through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$, it is necessary that $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$. We proceed to demonstrate that under this bound for $A$, there exists a $\gamma_{s}$ that is initially in $\mathfrak{S}_{A}(\sigma,\tau)$.
@@
We have
@@
For any parameter $s\geq 0$, the integral curve $\gamma_{s}$ initially enters $\mathcal{E}^{+}$ and $P_{X}>0$ holds. From Proposition 5.1 we know that $P_{X}+Q_{X}<0$ for each $k\in[0,1]$. Assume $\omega>0$ once $\gamma_{s}$ is away from $p_{0}^{+}$, and suppose $P\leq 0$ initially. The inequality $Q+P>0$ holds initially by (4.16). As $d_{1}\geq 2$, we have
@@
near $p_{0}^{+}$. Hence, the derivative $P^{\prime}$ in (4.3) is initially positive, which is a contradiction. The proof of the case $\omega<0$ is almost verbatim. ∎
@@
Let $s_{\bullet}=\frac{d_{1}}{\tau}=\frac{d_{1}\sigma}{(d_{2}-1)\sigma-(d_{1}+1)}$. The integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(k,l)$ if $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$.
@@
By (3.1), we have
By the monotonicity of $S$ and $T$, the inequalities $S\leq\sigma$ and $T\geq\tau$ hold along $\gamma_{s_{\bullet}}$ once the integral curve leaves $p_{0}^{+}$. Since $\gamma_{s_{\bullet}}$ is initially tangent to $\{S=\sigma\}$, the $(k,l)$-projection of $\gamma_{s_{\bullet}}$ is tangent to the level curve $\omega(A,k,l)=0$ at $(0,0)$ along the vector $v=(1,\sigma)$. By (4.15), we have $\left.\mathrm{Hess}(\omega)\right|_{(0,0)}(v,v)>0$ given that $A<\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$. By Proposition 4.6, the inequality $P>0$ holds initially along $\gamma_{s_{\bullet}}$. Therefore, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$. ∎
Proposition 4.7 provides a condition on $A$ to form a barrier for $\gamma_{s_{\bullet}}$ locally, which means that $\gamma_{s_{\bullet}}$ may still escape $\mathfrak{S}_{A}(\sigma,\tau)$ through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ eventually. In the following, we explore the condition on $A$ to make $\mathfrak{S}_{A}(\sigma,\tau)$ a global barrier so that $\gamma_{s_{\bullet}}$ can only escape through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}>0\}$.
@@
For any $d_{2}\geq d_{1}\geq 2$, the inequalities $\omega_{2}<0$ and $\omega_{0}>0$ hold for any $k\in[0,1]$.
@@
In the following, we specify how $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ gives rise to a closed interval for $l$. In terms of $k$ and $l$, the inequality $S\leq\sigma$ is equivalent to
@@
The coefficients $\beta_{0}$ and $\beta_{1}$ are positive for each $k\in[0,1]$ by Proposition 5.2. Furthermore, from Proposition 5.3 we know that
@@
Denote the smaller positive root of $\zeta(A,k,l)$ as $\nu_{2}(A,k)$. Proposition 4.9 guarantees the existence of the closed interval $[\nu_{2}(A,k),\nu_{1}(A,k)]$ for any $(A,k)\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)\times(0,1)$.
@@
Define $A_{1}=\min\limits_{k\in[0,1]}\left\{\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\frac{\beta_{2}}{\omega_{0}}\right\}$, where
@@
(4.24) $\begin{split}&\omega\left(A,k,\nu_{1}(A,k)\right)=\frac{(d_{1}-1)k^{2}}{d_{1}^{2}(n-1)}\left(\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\frac{\beta_{2}}{A}-\omega_{0}\right)\\ \end{split}$
By Proposition 5.4 in the Appendix, the factor $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\beta_{2}$ is positive for any $k\in[0,1]$. As $\omega_{0}$ is also positive by Proposition 4.8, we know that $A_{1}>0$. The proof is complete. ∎
@@
Therefore, the inclusion $[\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ holds with any sufficiently small $k>0$. Therefore, the inclusion can be extended to all $k\in(0,1)$ if $\nu_{2}(A,k)>\mu_{2}(A,k)$ holds for all $k\in(0,1)$. Such a condition puts another restriction on the upper bound for $A$.
@@
with coefficients $\theta_{i}(k)$ be polynomials; see Proposition 5.5. Furthermore, by Proposition 5.6, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ for any $k\in[0,1]$. The coefficient $\theta_{0}$ only vanishes at $k=0$. We also have
@@
Since $\chi_{d_{1},d_{2}}\leq\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$ by Proposition 4.7. Since $\chi_{d_{1},d_{2}}\leq A_{1},A_{2}$, the interval $[\nu_{2}(A,k),\nu_{1}(A,k)]$ exists and $[\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ for any $k\in(0,1)$ by Proposition 4.9, Proposition 4.10, and Proposition 4.11. Therefore, the inequality $Q>0$ holds on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}\cap\{X_{1}>X_{2}>0\}$. By (4.3) and Proposition 4.3, the integral curve $\gamma_{s_{\bullet}}$ intersects $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}\geq 0\}$. ∎
Lemma 4.2 and Lemma 4.12 proves Theorem 1.1.
5. Appendix
Proposition 5.1.

### v5

@@ 4. Global Analysis
Motivated by the derivative formula (4.5)-(4.7) and Proposition 4.6, we next analyze the sign of $Q$ on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. This is crucial in ruling out escaping through $\{P=0\}$ for suitable ranges of $A$, and also helps clarify whether $\gamma_{s_{\bullet}}$ lies in $\mathfrak{S}_{A}(\sigma,\tau)$ once it leaves $p_{0}^{+}$.
The set $\mathfrak{S}_{A}(\sigma,\tau)$ is equivalently defined as
@@
For any $d_{1},d_{2}\geq 2$, the inequality $P_{X}\geq 0$ holds on $[0,1]$ and only vanishes at $k=1$.
@@
The coefficient of each $k^{i}$ is non-neagative on $[0,1]$ since $d_{1},d_{2}\geq 2$. Therefore, the polynomial is positive on $[0,1]$. ∎
@@
The critical point $p_{0}^{+}$, whose $(k,l)$-coordinate is $(0,0)$, is on the boundary of $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$. If $A=0$, the product Einstein metric is represented by the algebraic curve $\mathcal{E}\cap\{X_{2}=0\}\cap\left\{R_{2}-\frac{1-H^{2}}{n}=0\right\}$, at which both $P$ and $Q$ identically vanish. This is reflected by $\omega(0,0,l)$ being identically zero.
For any $A\geq 0$, both $\omega(A,k,l)$ and its gradient vanish at $(0,0)$. To determine the sign of $\omega(A,k,l)$ near $p_{0}^{+}$, we compute the Hessian of the function at $(0,0)$,
@@
By the second derivative test, we have the following proposition.
@@
Note that $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$ is not larger than the upper bound in Proposition 4.6 provided $d_{1},d_{2}\geq 2$. In particular, Propositions 4.6 and 4.8, together with the derivative formula (4.7), show that a necessary condition for $Q\geq 0$ on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ is $A\in\left[0,\tfrac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right).$
From now on we restrict to this range of $A$. To prove that $\gamma_{s_{\bullet}}$ indeed enters $\mathfrak{S}_{A}(\sigma,\tau)$, we examine in turn the defining inequalities of $\mathfrak{S}_{A}(\sigma,\tau)$. The bounds involving $S\leq\sigma$ and $T\geq\tau$ are immediate from the linearized solution near $p_{0}^{+}$ and the monotonicity of these functions, as shown in the following proposition.
Proposition 4.9.
Along the integral curve $\gamma_{s_{\bullet}}$ one has $T\geq\tau$ once the curve leaves $p_{0}^{+}$. One also has $S\leq\sigma$ initially provided that $P>0$ initially.
Proof.
By (3.2), we have
By definitions, we have $S,T>0$ in $(\mathcal{E}^{+})^{\circ}$. Moreover, by (4.6), the function $T$ is monotonic in $(\mathcal{E}^{+})^{\circ}$. By (4.5), the function $S$ is monotonic provided that $P>0$. Therefore, the inequalities $S\leq\sigma$ and $T\geq\tau$ hold along $\gamma_{s_{\bullet}}$ once the integral curve leaves $p_{0}^{+}$, given that $P>0$ initially. ∎
Remark 4.10.
The rational functions $S$ and $T$ are not defined at $p_{0}^{+}$ since $X_{2}=Z=0$ at $p_{0}^{+}$. Nevertheless, the conditions $S=\sigma$ and $T=\tau$ extend smoothly across $p_{0}^{+}$ via the relation
$\tilde{S}:=Z\left(X_{1}+\frac{d_{2}}{2d_{1}}X_{2}\right)-\sigma YX_{2}=0,\quad\tilde{T}:=\frac{1-H^{2}}{n}-\tau YZ=0.$
Consider the unstable eigenvector $s_{\bullet}v_{1}+v_{2}$, along which $\gamma_{s_{\bullet}}$ emanates $p_{0}^{+}$. Since
$\nabla\tilde{S}(p_{0}^{+})=\begin{bmatrix}0\\ -\frac{\sigma}{d_{1}}\\ 0\\ \frac{1}{d_{1}}\end{bmatrix},\quad\nabla\tilde{T}(p_{0}^{+})=\begin{bmatrix}-\frac{2d_{1}}{n}\\ -\frac{2d_{2}}{n}\\ 0\\ -\frac{\tau}{d_{1}}\end{bmatrix}$,
we have
$\langle\nabla\tilde{S}(p_{0}^{+}),s_{\bullet}v_{1}+v_{2}\rangle=\langle\nabla\tilde{T}(p_{0}^{+}),s_{\bullet}v_{1}+v_{2}\rangle=0.$
Under this smooth extension, the unstable eigenvector is tangent to the hypersurfaces $\{S=\sigma\}$ and $\{T=\tau\}$ at $p_{0}^{+}$. We work with $S$ and $T$ (instead of $\tilde{S}$ and $\tilde{T}$) because their derivatives along integral curves, (4.5)–(4.6), are simpler and allow one to determine the initial sign of $S-\sigma$ and $T-\tau$ along $\gamma_{s_{\bullet}}$. In contrast, the linearized solution provides no information about initial signs of $\tilde{S}$ and $\tilde{T}$, so they are less effective for our purpose.
It is left to show that $P>0$ initially along $\gamma_{s_{\bullet}}$. Since
$\nabla P(p_{0}^{+})=\begin{bmatrix}\frac{2}{n}\\ \frac{d_{1}(d_{2}-d_{1})-n}{nd_{1}^{2}}\\ 0\\ \frac{d_{2}-1}{d_{1}^{2}}\end{bmatrix}$
is perpendicular to both $v_{1}$ and $v_{2}$, the initial sign of $P$ along $\gamma_{s}$ cannot be read off from the linearized solution (3.2) (similar to the situation for $\tilde{S}$ and $\tilde{T}$ in Remark 4.10). To resolve this, we consider the auxiliary function $\omega$, whose sign controls the initial behavior of $P$, as shown in the next proposition.
@@
Recall that the $(k,l)$-coordinate of $p_{0}^{+}$ is $(0,0)$, and we have
$P_{X}(0)=\frac{d_{1}(d_{1}-1)}{n-1}>0.$
Hence for any $\gamma_{s}$ with $s>0$, the integral curve initially satisfies $P_{X}>0$ in a small neighborhood of $p_{0}^{+}$. We have
@@
By Proposition B.1 we know that $P_{X}+Q_{X}<0$ for each $k\in[0,1]$.
Consider first the case $\omega>0$ once $\gamma_{s}$ leaves $p_{0}^{+}$. If we assume $P\leq 0$ initially, then the two terms on the right-hand side of (4.22) are both non-negative: the first term is strictly positive (since $\omega>0$, $P_{X}>0$ and $Y^{2}X_{1}>0$ near $p_{0}^{+}$), while the second term is non-negative because $\frac{P_{X}+Q_{X}}{P_{X}}<0$ and $P\leq 0$. Thus $Q+P>0$ initially.
Next we examine the sign of $P^{\prime}$ near $p_{0}^{+}$. We claim that the first term in (4.7) is non-negative, since we assume $P\leq 0$ and
@@
near $p_{0}^{+}$. Together with $Q+P>0$ and $X_{1}-X_{2}=\frac{1}{d_{1}}>0$ initially (which makes the second term in (4.7) positive), this implies that $P^{\prime}$ is initially strictly positive, contradicting the assumption $P\leq 0$ initially.
The argument for the case $\omega<0$ is analogous: if $P\geq 0$ initially then by (4.22) we obtain $Q+P<0$ initially. The inequality (4.23) still holds. By the assumption that $P\geq 0$, the first term in (4.7) is non-positive. Since $X_{1}-X_{2}=\frac{1}{d_{1}}$ and $Q+P<0$ initially, the second term in (4.7) is negative. Hence, the derivative $P^{\prime}$ is negative initially, a contradiction.
Therefore, whenever $\omega$ is not identically zero along $\gamma_{s}$, the sign of $\omega$ agrees with the sign of $P$ once the integral curve leaves $p_{0}^{+}$. ∎
@@
The integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(k,l)$ if $A\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)$.
@@
In terms of $(k,l)$, the relation $S=\sigma$ is equivalent to
$l=\sigma\frac{2d_{1}k}{2d_{1}+d_{2}k}.$
By Remark 4.10, the integral curve $\gamma_{s_{\bullet}}$ is tangent to $\{S=\sigma\}$ after smooth extension at $p_{0}^{+}$. Therefore, the $(k,l)$-projection of $\gamma_{s_{\bullet}}$ is tangent to $v=(1,\sigma)$ at $(0,0)$. Recall that $\omega$ and its gradient vanish at $(0,0)$. By (4.21), we have
$\left.\mathrm{Hess}(\omega)\right|_{(0,0)}(v,v)=\frac{2(d_{1}-1)}{n-1}\left(\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}}\frac{1}{A}-(d_{1}d_{2}-d_{2}+4)\right)>0$
given that $A<\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$. Hence, the function $\omega$ is positive initially along $v$. By Proposition 4.11, the inequality $P>0$ holds initially along $\gamma_{s_{\bullet}}$. Together with Proposition 4.9, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$. ∎
Proposition 4.12 provides a condition on $A$ to form a barrier for $\gamma_{s_{\bullet}}$ locally, which means that $\gamma_{s_{\bullet}}$ may still escape $\mathfrak{S}_{A}(\sigma,\tau)$ through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ eventually. In the following, we explore the condition on $A$ to preserve the positivity of $P$ (or equivalently the positivity of $\omega$) when $\gamma_{s_{\bullet}}$ is in $\mathfrak{S}_{A}(\sigma,\tau)$, making $\mathfrak{S}_{A}(\sigma,\tau)$ a global barrier so that $\gamma_{s_{\bullet}}$ can only escape through $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}>0\}$.
@@
For any $d_{1},d_{2}\geq 2$, the inequalities $\omega_{2}<0$ and $\omega_{0}>0$ hold for any $k\in[0,1]$.
@@
In the following, we specify how $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}$ gives rise to a closed interval for $l$. In terms of $(k,l)$, the inequality $S\leq\sigma$ is equivalent to
@@
The coefficients $\beta_{0}$ and $\beta_{1}$ are positive for each $k\in[0,1]$ by Proposition B.2. Furthermore, from Proposition B.3 we know that
@@
Denote the smaller positive root of $\zeta(A,k,l)$ as $\nu_{2}(A,k)$. Proposition 4.14 guarantees the existence of the closed interval $\nu_{2}(A,k),\nu_{1}(A,k)]$ for any $(A,k)\in\left(0,\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}\right)\times(0,1)$.
@@
Define $A_{1}=\min\limits_{k\in[0,1]}F$, where $F:=\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\frac{\beta_{2}}{\omega_{0}}$, and
@@
[…]
By Proposition B.4 in the Appendix, the factor $\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(2d_{1}+d_{2}k)}\beta_{2}$ is positive for any $k\in[0,1]$. As $\omega_{0}$ is also positive by Proposition 4.13, we know that $A_{1}>0$. The proof is complete. ∎
@@
Therefore, the inclusion $\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ holds with any sufficiently small $k>0$. Therefore, the inclusion can be extended to all $k\in(0,1)$ if $\nu_{2}(A,k)>\mu_{2}(A,k)$ holds for all $k\in(0,1)$. Such a condition puts another restriction on the upper bound for $A$.
@@
with coefficients $\theta_{i}(k)$ be polynomials; see Proposition B.5. Furthermore, by Proposition B.6, we have $\theta_{2}<0$ and $\theta_{0}\geq 0$ for any $k\in[0,1]$. The coefficient $\theta_{0}$ only vanishes at $k=0$. We also have
@@
Since $\chi_{d_{1},d_{2}}\leq\frac{d_{2}(d_{2}-1)^{2}}{d_{1}^{2}(d_{1}d_{2}-d_{2}+4)}$, the integral curve $\gamma_{s_{\bullet}}$ is initially in $\mathfrak{S}_{A}(\sigma,\tau)$ by Proposition 4.12. Since $\chi_{d_{1},d_{2}}\leq A_{1},A_{2}$, the interval $\nu_{2}(A,k),\nu_{1}(A,k)]$ exists and $\nu_{2}(A,k),\nu_{1}(A,k)]\subset(\mu_{2}(A,k),\mu_{1}(A,k))$ for any $k\in(0,1)$ by Proposition 4.14, Proposition 4.15, and Proposition 4.16. Therefore, the inequality $Q>0$ holds on $\mathfrak{S}_{A}(\sigma,\tau)\cap\{P=0\}\cap\{X_{1}>X_{2}>0\}$. By (4.5)-(4.6) and Proposition 4.6, the integral curve $\gamma_{s_{\bullet}}$ intersects $\mathfrak{S}_{A}(\sigma,\tau)\cap\{X_{1}=X_{2}\geq 0\}$. ∎
Lemma 4.4 and Lemma 4.17 prove Theorem 1.1.
Appendix A Ricci–flat metrics with first integrals
Theorem A.1.
If $(d_{1},d_{2},A)$ satisfies
(A.1) $A=\frac{d_{2}(d_{2}-1)^{2}}{4d_{1}^{2}(d_{1}+1)^{2}}(d_{2}+4d_{1}-d_{1}d_{2}),$
the set $\mathcal{L}$ is invariant. The Ricci-flat metric represented by $\gamma_{\infty}$ is integrable.
Proof.
With $H=1$ in $\mathcal{RF}^{+}$, we have
(A.2) $\begin{split}&\left.\left(X_{1}-Y+\frac{d_{2}}{2d_{1}}\frac{d_{2}-1}{d_{1}+1}Z\right)^{\prime}\right|_{\mathcal{L}}\\ &=X_{1}(G-1)+(d_{1}-1)Y^{2}+AZ^{2}-Y(G-X_{1})+\frac{d_{2}}{2d_{1}}\frac{d_{2}-1}{d_{1}+1}Z(G+X_{1}-2X_{2})\\ &=X_{1}(G-d_{1}X_{1}-d_{2}X_{2})+(d_{1}-1)Y^{2}+AZ^{2}-Y(G-X_{1})+\frac{d_{2}}{2d_{1}}\frac{d_{2}-1}{d_{1}+1}Z(G+X_{1}-2X_{2})\end{split}$
and
(A.3) $\begin{split}&\left.\left(X_{2}-\frac{d_{2}-1}{d_{1}+1}Z\right)^{\prime}\right|_{\mathcal{L}}\\ &=X_{2}(G-1)+(d_{2}-1)YZ-\frac{2d_{1}}{d_{2}}AZ^{2}-\frac{d_{2}-1}{d_{1}+1}Z(G+X_{1}-2X_{2})\\ &=X_{2}(G-d_{1}X_{1}-d_{2}X_{2})+(d_{2}-1)YZ-\frac{2d_{1}}{d_{2}}AZ^{2}-\frac{d_{2}-1}{d_{1}+1}Z(G+X_{1}-2X_{2}).\end{split}$
Replace $(X_{1},X_{2})$ with the linear expressions in $(Y,Z)$, the above equations respectively become
$\left(A-\frac{d_{2}(d_{2}-1)^{2}}{4d_{1}^{2}(d_{1}+1)^{2}}(d_{2}+4d_{1}-d_{1}d_{2})\right)Z^{2},\quad\left(\frac{1}{2d_{1}}\frac{(d_{2}-1)^{2}}{(d_{1}+1)^{2}}(d_{2}+4d_{1}-d_{1}d_{2})-\frac{2d_{1}}{d_{2}}A\right)Z^{2},$
which vanish if (A.1) holds. ∎
$(d_{1},d_{2},A)$ Principal Orbit/Hypersurface Holonomy Source
$\left(1,d,\frac{d(d-1)^{2}}{4}\right)$ $\mathbb{S}^{1}$-bundle associated to $\mathcal{O}(d+2)$ Kähler [BB82]
$\left(2,4,1\right)$ Twistor bundle over $\mathbb{HP}^{1}$ $G_{2}$ [BS89]
$\left(3,3,\frac{1}{8}\right)$ $\mathbb{S}^{3}\times\mathbb{S}^{3}$
$\left(3,4,\frac{1}{4}\right)$ Hopf fibration over $\mathbb{HP}^{1}$ $\mathrm{Spin}(7)$
$(2,8,0)$ Product space Generic [DW99]
$(3,6,0)$
$(5,5,0)$
Table 1. Examples that satisfy (A.1)
Appendix B Computations of Rational functions
Proposition B.1.